OpenStudy (anonymous):

Find the absolute maximum and absolute minimum values of the function f (x) = x^4−10x^2−7 on each of the indicated intervals. Enter None for any absolute extrema that do not exist. (A) Interval = [−3,−1]. Absolute maximum =-16 Absolute minimum = (B) Interval = [−4,1]. Absolute maximum =89 Absolute minimum = (C) Interval = [−3,4]. Absolute maximum =89 Absolute minimum = I know the absolute value for all would be the same, still trying to find it...any help?

OpenStudy (anonymous):

Absolute minimum I meant!

OpenStudy (anonymous):

this is three different questions right?

OpenStudy (anonymous):

Yeah...but the absolute minimum...i believe is the same for all

OpenStudy (anonymous):

\[f (x) = x^4−10x^2−7\] \[f'(x)=4x^3-20x=4x(x^2-5)=4x(x+\sqrt{5})(x-\sqrt{5})\]

OpenStudy (anonymous):

is that right so far? critical points at \(x=0,x=-\sqrt{5},x=\sqrt{5}\)

OpenStudy (anonymous):

so \(-\sqrt{5}, \sqrt{5}\) will give a local min, and 0 will be a local max. now you only have to worry about what interval your are in

OpenStudy (anonymous):

So far, I'm not sure...with those answers...I did it that way too but i didn't quite get it...

OpenStudy (anonymous):

Hmm is another ex....with all correct answers Find the absolute maximum and absolute minimum values of the function f (x) = x4−4x2−2 on each of the indicated intervals. Enter None for any absolute extrema that do not exist. (A) Interval = [−3,−1]. Absolute maximum =43 Absolute minimum =-6 (B) Interval = [−4,1]. Absolute maximum =190 Absolute minimum =-6 (C) Interval = [−3,4]. Absolute maximum =190 Absolute minimum =-6 Is the finding the roots work here as well?